OpenStudy (anonymous):

Rewrite the following equation with only sinx and cosx. sin2x-cos2x

OpenStudy (ajprincess):

sin2x=2sinxcosx cos2x=cos^2x-sin^2x =2cos^2x-1 =1-2sin^2x Can u do it @ballerinaxbb?

OpenStudy (anonymous):

Got it, 2sinxcosx-1 + 2sin^2x

OpenStudy (anonymous):

Thank you!

OpenStudy (ajprincess):

welcome:)

OpenStudy (anonymous):

I posted another similar question, do you think you can help?